# FactCheXcker: Mitigating Measurement Hallucinations in Chest X-ray Report Generation Models

### v1

@@ Abstract
Medical vision-language model models often struggle with generating accurate quantitative measurements in radiology reports, leading to hallucinations that undermine clinical reliability. We introduce FactCheXcker, a modular framework that de-hallucinates radiology report measurements by leveraging an improved query-code-update paradigm. Specifically, FactCheXcker employs specialized modules and the code generation capabilities of large language models to solve measurement queries generated based on the original report. After extracting measurable findings, the results are incorporated into an updated report. We evaluate FactCheXcker on endotracheal tube placement, which accounts for an average of 78% of report measurements, using the MIMIC-CXR dataset and 11 medical report-generation models. Our results show that FactCheXcker significantly reduces hallucinations, improves measurement precision, and maintains the quality of the original reports. Specifically, FactCheXcker improves the performance of all 11 models and achieves an average improvement of 94.0% in reducing measurement hallucinations measured by mean absolute error.
@@ 1 Introduction
Figure 2: Distribution of measurable findings across different models. The radial axis shows the percentage of each finding type relative to the total measurable findings for each model. Notably, endotracheal tubes (shown in red) represent, on average, 78% of the measurements across all models. The remaining categories account for an average of 12.1%, 8.4%, 0.7%, and 0.5% of measurements, respectively.
Motivated by this, we present FactCheXcker. This extensible tool-based pipeline reevaluates and updates measurements in model-generated radiology reports without retraining or modifying the original model. FactCheXcker follows a query-code-update paradigm. Specifically, the framework automatically generates measurement queries based on the original radiology report, such as “measure [XYZ]”. It generates executable code based on a domain-specific API that solves the measurement query and returns a result. Furthermore, we extend the pipeline to incorporate the results into an updated report automatically. Thus, the FactCheXcker pipeline can iteratively improve upon any radiology report without retraining the original model. The flexible format allows FactCheXcker to solve a wide range of measurement queries by combining modular methods to perform more complex tasks, such as finding the distance between two anatomies or counting the number of nodules.
Table 1: Examples of specific measurements of various objects in medical reports.
@@
Our investigation begins with analyzing the most common measurable findings mentioned in radiology reports across 11 report generation models and real-world data. Figure 2 shows that endotracheal tubes (ETT) comprise most concrete measurements. ETTs assist patients with mechanical ventilation but require regular monitoring through chest X-rays to ensure proper positioning, thereby minimizing the risk of faulty intubations. Accurate placement is critical, as an improperly positioned ETT can result in complications such as hypoxia, pneumothorax, or even death [12]. Radiologists typically measure the distance between the carina and the tip of the ETT in centimeters, recommending repositioning if the distance is outside the optimal range. Precise measurement is, therefore, crucial to ensuring proper patient care. See Table 1 for examples of measurements in radiology reports.
To evaluate the effectiveness of FactCheXcker, we apply it to the task of assessing ETT placement. We evaluated FactCheXcker on the MIMIC-CXR dataset [17] for the ETT placement task using 11 different report-generation models. Our results demonstrate that FactCheXcker significantly reduces measurement hallucinations and improves ETT detection and placement precision, all while preserving the overall quality of the original reports. In summary, our main contributions are as follows:
@@
  • We show that applying FactCheXcker improves the performance of all models and significantly reduces hallucinated ETT placement measurements, achieving an average improvement of 94% in mean absolute error.
@@ 2 Related Work
Medical Report Generation Models. Medical report generation has seen significant advances with the emergence of large language models. Early approaches utilized CNN-RNN architectures to generate reports from medical images [35, 44]. At the same time, more recent works have adopted transformer-based architectures for improved performance [4, 39, 19, 40]. These models typically combine vision encoders, such as Vision Transformer [9] and Swin Transformer [24], with text decoders to generate detailed clinical descriptions. For instance, CheXpertPlus baseline models [3] integrates Swinv2 [25] architecture with a two-layer BERT decoder [18] for medical report generation. In addition, more versatile foundation models have emerged, such as MedVersa [48] which coordinates multimodal inputs and outputs for varying tasks, RadFM [42] which supports both 2D and 3D scans with visual-language interleaving, and CheXagent [5] which takes an instruction-tuned approach with a specialized vision encoder bridged with language modalities. Recently, MAIRA-2 [2] combines radiology-specific image encoders with LLMs for grounded report generation. Despite these architectural advancements, a critical limitation remains: these models often struggle with accurately interpreting fine-grained quantitative information and spatial relationships within medical images, leading to inaccuracies in critical measurements, such as endotracheal tube placement. This highlights the need for specialized approaches to handle quantitative measurements in medical imaging.
@@
Enhancing LLM Capabilities with External APIs. In recent years, integrating large language models with external APIs has become a popular approach for expanding action spaces and tackling complex tasks [33, 22, 20]. For instance, HuggingGPT [34] utilizes models from the HuggingFace platform to dynamically call specialized models and address diverse user needs. Similarly, ViperGPT [38] uses Codex, an LLM designed to convert text descriptions into executable Python code, allowing the model to complete tasks by directly running the generated code. Inspired by these works, we propose leveraging LLMs and APIs to mitigate measurement hallucination.
@@
Fact-Checking and Mitigating Hallucination in LLMs. Hallucinations represent a significant challenge in LLMs, where the models generate content that is not grounded in the input signal but instead relies on their internal knowledge [16]. This issue has increased focus on fact-checking and mitigating hallucinations across multiple modalities [11, 23, 1]. One line of research to address this problem has focused on hallucination detection frameworks and fact-checking mechanisms for LLM outputs. FacTool [6] proposes a task- and domain-agnostic framework that leverages GPT for extracting and verifying claims. ProgramFC [30] and FActScore [28] both approached complex fact-checking by breaking down statements into simpler subtasks or atomic statements that could be individually verified against a knowledge source. In multimodal LLMs, approaches [21, 27, 41] have been proposed that include using binary questions to assess object presence in visual inputs or leveraging Visual Question Answering (VQA) methods to generate negative statements indicating the absence of particular objects in visual scenes. Despite progress, a significant gap remains in the medical domain for real-life clinical usage. The precise measurement requirements of medical imaging, such as identifying the size of lesions, or the exact positioning of medical devices, necessitate specific methods to mitigate the hallucination of such quantitative aspects.
@@ 3 Method
In this section, we introduce FactCheXcker, an extensible and modularized pipeline designed to eliminate hallucinated measurements in radiology reports. As illustrated in Figure 1, FactCheXcker consists of three main components: Query Generator, Code Generator, and Report Updater. When provided with a medical image and its corresponding AI-generated report that may contain hallucinated measurements, the Query Generator identifies potential measurement discrepancies in the report, the Code Generator creates and executes specialized code to obtain precise measurements from the image, and the Report Updater integrates these verified measurements into the report.
@@ 3.1 Task Definition
We categorize hallucinations in medical report generation models into two types: object hallucination and measurement hallucination. Object hallucination occurs when the model erroneously predicts the presence of specific diseases in particular anatomical structures, such as incorrectly identifying pneumonia in the right lung lobe. On the other hand, measurement hallucination refers to inaccurate predictions of numerical values related to quantifiable findings, such as the size of lesions or distances between devices and anatomical structures. The extensible design of FactCheXcker allows it to address both categories of hallucinations. However, in this paper, we primarily focus on detecting and correcting measurement hallucinations in AI-generated radiology reports, as this issue is more prevalent and critically requires our approach.
@@ 3.2 FactCheXcker Framework
Query Generator. This module ($m_{query}$) receives the original report and infers what information needs to be re-measured and verified. It then generates a list of measurement queries that guide subsequent steps. For instance, if the original report was “The endotracheal tube terminates 2.3 cm above the carina. There is a new dense right central opacity measuring about 6 cm x 3 cm”; the framework would produce the queries ”measure the distance between the endotracheal tube and the carina” and ”measure the dimensions of the right central opacity”. For this purpose, we utilize GPT-4o mini [14] as the query generator, and details on the prompt used can be found in the Appendix. These queries are then passed on to the next module, the Code Generator.
Code Generator. This module ($m_{code}$) receives the measurement query and a high-level API description. It dynamically generates Python code to answer the query based on the methods provided by the API. We have designed a comprehensive toolbox for identifying and correcting measurement hallucinations in radiology reports, as detailed in the following:
@@
Report Updater. This module ($m_{update}$) takes the original report together with all the results of the measurement queries to update the report. Specifically, it updates a sentence if a new value is presented, or removes a phrase entirely if the object was hallucinated. Moreover, since the old interpretation may be outdated in light of new evidence, the module references measurement guidelines for each result to determine if the ETT placement should be updated. The pipeline concludes with returning the updated report. For instance, in the case of ETT placement, where the ideal position is 5 ± 2 cm above the carina  [10], if the updated measurement falls outside this range, the Report Updater will flag this discrepancy and update the report to indicate the need for repositioning. We notice that the numeric reasoning of GPT-4o mini [14] is currently insufficient to reliably extract the inference. For instance, the language model fails to identify if a measurement is below or above a threshold value. Therefore, in this version of the pipeline, we implement a strict rule-based approach to infer the placement category of the endotracheal tube. If the distance is between 3-7 centimeters, the placement is deemed as “correct”. Otherwise, the placement is updated to “incorrect”.
@@ 4.2 Tool Modules
  • find(“carina”, “endotracheal tube”): We fine-tune CarinaNet [29] on a diverse private dataset from 22 hospitals. The CarinaNet module, which we call CarinaNet+, outputs an (x, y)-coordinate for both the carina and the endotracheal tube along with confidence scores. See Table 2 for the performance on the MIMIC-CXR test set.
@@ 4.3 Baselines and Implementation Details
Baseline Models. We demonstrate FactCheXcker’s effectiveness across 11 different medical report generation models, including CheXagent [5], CheXpertPlus [3], GPT-4V [45], LLM-CXR [19], MAIRA-2 [2], MedVersa [48], RadFM [42], RaDialog [31], RGRG [39], VLCI [4]. For all models, we follow the official protocols to generate reports on the MIMIC-CXR test set. Detailed descriptions and implementation specifications are provided in the Appendix.
Measurement Extraction. We implement a systematic extraction process to extract labels of the concrete measurements made in the ground truth- and model-generated reports. To optimize API usage, we first filter reports containing specific measurements using keywords: “cm”, “mm”, “centimeter(s)”, “millimeter(s)”, “measure(s)”. We then utilize GPT-4o mini to extract object names with associated concrete measurements from both ground truth and model-generated reports. The detailed extraction prompt is provided in the Appendix. We categorize the identified measurements into five main categories: Endotracheal tube, Other tubes/catheters (e.g., PICC, central venous line), Lesion (e.g., opacity, mass, nodule), Pneumothorax, and Others (e.g., calcification, balloon pump). Notably, endotracheal tubes account for the most significant proportion of measurements, with an average mention percentage of 78.3% across all models.
Endotracheal Tube Annotation. After extraction of the measurements from the MIMIC-test set, there remained 45 reports where the ground truth report mentioned an endotracheal tube as present, but did not give a concrete measurement. To give more opportunity for comparisons, a trained radiologist annotated the location of the carina and the endotracheal tube tip in the 45 images. Then, we computed the distance between the centers of the annotated bounding boxes as the distance between the carina and the endotracheal tube. This result was passed to the update report-module of our framework, and thus the result was “injected” into the ground truth reports for later comparisons.
@@
ETT Presence ETT Measurement ETT Placement
@@
CheXpertPlus 0.64 0.68 2.47 1.30 90.0% 3.58 1.81 98.0% 0.64 0.73
GPT4V 0.20 0.58 3.56 1.15 210.0% 16.95 3.59 372.0% 0.77 0.85
MAIRA-2 0.63 0.68 1.63 1.36 20.0% 2.30 1.84 25.0% 0.62 0.72
CheXagent 0.70 0.76 2.58 1.20 115.0% 5.38 2.45 120.0% 0.74 0.81
RaDialog 0.67 0.69 1.23 1.20 3.0% 1.78 1.71 4.0% 0.67 0.77
Cvt2distilgpt2 0.71 0.72 3.78 1.35 180.0% 6.30 2.21 185.0% 0.59 0.70
MedVersa 0.73 0.80 1.46 1.23 19.0% 2.03 1.62 25.0% 0.71 0.77
LLM-CXR 0.08 0.58 2.80 1.25 124.0% 21.54 2.27 849.0% 0.62 0.79
RadFM 0.28 0.47 1.69 1.10 54.0% 21.12 12.22 73.0% 0.67 0.83
VLCI 0.25 0.54 3.69 1.67 121.0% 23.06 8.35 176.0% 0.70 0.75
RGRG 0.50 0.60 2.38 1.22 95.0% 3.78 1.74 117.0% 0.64 0.74
Average 0.49 0.65 2.48 1.28 94.0% 9.80 3.62 171.0% 0.67 0.77

### v2

@@ Abstract
Medical vision-language models often struggle with generating accurate quantitative measurements in radiology reports, leading to hallucinations that undermine clinical reliability. We introduce FactCheXcker, a modular framework that de-hallucinates radiology report measurements by leveraging an improved query-code-update paradigm. Specifically, FactCheXcker employs specialized modules and the code generation capabilities of large language models to solve measurement queries generated based on the original report. After extracting measurable findings, the results are incorporated into an updated report. We evaluate FactCheXcker on endotracheal tube placement, which accounts for an average of 78% of report measurements, using the MIMIC-CXR dataset and 11 medical report-generation models. Our results show that FactCheXcker significantly reduces hallucinations, improves measurement precision, and maintains the quality of the original reports. Specifically, FactCheXcker improves the performance of all 11 models and achieves an average improvement of 135.0% in reducing measurement hallucinations measured by mean absolute error. Code is available at https://github.com/rajpurkarlab/FactCheXcker.
Equal contribution.   Corresponding author.
@@ 1 Introduction
Figure 2: Distribution of measurable findings across different models. The radial axis shows the percentage of each finding type relative to the total measurable findings for each model. Notably, endotracheal tubes (shown in red) represent, on average, 78% of measurements across all models. The remaining categories account for an average of 12.1%, 8.4%, 0.7%, and 0.5% of measurements, respectively.
Motivated by this, we present FactCheXcker. This extensible tool-based pipeline reevaluates and updates measurements in model-generated radiology reports without retraining or modifying the original model. FactCheXcker follows a query-code-update paradigm. Specifically, the framework automatically generates measurement queries based on the original radiology report, such as “measure [XYZ]”. It generates executable code based on a domain-specific API that solves the measurement query and returns a result. Furthermore, we extend the pipeline to automatically incorporate the results into an updated report. Thus, the FactCheXcker pipeline can iteratively improve any radiology report without retraining the original model. The flexible format allows FactCheXcker to solve a wide range of measurement queries by combining modular methods to perform more complex tasks, such as finding the distance between two anatomies or counting the number of nodules.
Table 1: Examples of specific measurements of various objects in radiology reports.
@@
Our investigation begins with analyzing the most common measurable findings mentioned in radiology reports across 11 report generation models and real-world data. Figure 2 shows that endotracheal tubes comprise most concrete measurements. ETTs assist patients with mechanical ventilation but require regular monitoring through chest X-rays to ensure proper positioning, thereby minimizing the risk of faulty intubations. Accurate placement is critical, as an improperly positioned ETT can result in complications such as hypoxia, pneumothorax, or even death [12]. Radiologists typically measure the distance between the carina and the tip of the ETT in centimeters and recommend repositioning if it is outside the optimal range. Precise measurement is, therefore, crucial to ensuring proper patient care. Table 1 lists measurements seen in radiology reports.
To evaluate the effectiveness of FactCheXcker, we apply it to the task of assessing ETT placement. We evaluate FactCheXcker on the MIMIC-CXR dataset [17] for the ETT placement task using 11 different report-generation models. Our results demonstrate that FactCheXcker significantly reduces measurement hallucinations and improves ETT detection and placement precision, all while preserving the overall quality of the original reports. In summary, our main contributions are as follows:
@@
  • We show that applying FactCheXcker improves the performance of all models and significantly reduces hallucinated ETT placement measurements, achieving an average improvement of 135% in mean absolute error.
@@ 2 Related Work
Medical Report Generation Models. Medical report generation has seen significant advances with the emergence of large language models. Early approaches utilized CNN-RNN architectures to generate reports from medical images [35, 44]. At the same time, more recent works have adopted transformer-based architectures for improved performance [4, 39, 19, 40]. These models typically combine vision encoders, such as Vision Transformer [9] and Swin Transformer [24], with text decoders to generate detailed clinical descriptions. For instance, CheXpertPlus baseline models [3] integrates Swinv2 [25] architecture with a two-layer BERT decoder [18] for medical report generation. In addition, more versatile foundation models have emerged, such as MedVersa [48] which coordinates multimodal inputs and outputs for varying tasks, RadFM [42] which supports both 2D and 3D scans with visual-language interleaving, and CheXagent [5] which takes an instruction-tuned approach with a specialized vision encoder bridged with language modalities. Recently, MAIRA-2 [2] combines radiology-specific image encoders with large language models (LLMs) for grounded report generation. Despite these architectural advancements, critical limitations remain: the models often struggle with accurately interpreting fine-grained quantitative information and spatial relationships within medical images, highlighting the need for specialized approaches.
@@
Enhancing LLM Capabilities with External APIs. Integrating LLMs with external APIs has become popular for expanding action spaces and tackling complex tasks [33, 22, 20]. For instance, HuggingGPT [34] utilizes models from the HuggingFace platform to dynamically call specialized models and address diverse user needs. Similarly, ViperGPT [38] uses Codex, an LLM designed to convert text descriptions into executable Python code, allowing the model to complete tasks by directly running generated code.
@@
Fact-Checking and Mitigating Hallucination in LLMs. Hallucinations represent a significant challenge in LLMs, where the models generate content that is not grounded in the input signal but instead relies on their internal knowledge [16]. This issue has increased focus on fact-checking and mitigating hallucinations across multiple modalities [11, 23, 1]. One line of research to address this problem has focused on hallucination detection frameworks and fact-checking mechanisms for LLM outputs. FacTool [6] proposes a task- and domain-agnostic framework that leverages GPT for extracting and verifying claims. ProgramFC [30] and FActScore [28] both approached complex fact-checking by breaking down statements into simpler subtasks or atomic statements that could be individually verified against a knowledge source. In multimodal LLMs, approaches [21, 27, 41] have been proposed that include using binary questions to assess object presence in visual inputs or leveraging Visual Question Answering (VQA) methods to generate negative statements indicating the absence of particular objects in visual scenes.
Despite progress, a significant gap remains in the medical domain for real-life clinical usage. Medical imaging requires precise measurements, such as identifying the size of lesions or the exact positioning of medical devices. Inspired by these works, we propose leveraging LLMs and APIs to mitigate measurement hallucination. Our method uses specialized medical tools for accurate anatomical calculations, and automatically updates verified measurements in the report. This approach has two main advantages: no assumption of model weight access and portability to resource-limited and/or privacy-centered hospital settings. Small and locally deployable tool modules can serve multiple models while maintaining data compliance, avoiding the need for extensive model tuning or external data transfer.
@@ 3 Method
In this section, we introduce FactCheXcker [^1] https://github.com/rajpurkarlab/FactCheXcker, an extensible and modularized pipeline designed to eliminate hallucinated measurements in radiology reports. As illustrated in Figure 1, FactCheXcker consists of three main components: Query Generator, Code Generator, and Report Updater. When provided with a medical image and its corresponding model-generated report that may contain hallucinated measurements, the Query Generator identifies potential measurement discrepancies in the report, the Code Generator creates and executes specialized code to obtain precise measurements from the image, and the Report Updater integrates these verified measurements into the report.
@@ 3.1 Task Definition
We define hallucination as making an erroneous measurement of the presence of a tube when it does not exist or making the wrong measurement. We categorize hallucinations in medical report generation models into two types: object hallucination and measurement hallucination. Object hallucination occurs when the model erroneously predicts the presence of specific diseases in particular anatomical structures, such as incorrectly identifying pneumonia in the right lung lobe. On the other hand, measurement hallucination refers to inaccurate predictions of numerical values related to quantifiable findings, such as the size of lesions or distances between devices and anatomical structures. The extensible design of FactCheXcker allows it to address both categories of hallucinations. However, in this paper, we primarily focus on detecting and correcting measurement hallucinations in model-generated radiology reports, as this issue is more prevalent and critically requires our approach.
@@ 3.2 FactCheXcker Framework
Query Generator. This module receives the original report and infers what information needs to be re-measured and verified. It then generates a list of measurement queries that guide subsequent steps. For instance, if the original report was “The endotracheal tube terminates 2.3 cm above the carina. There is a new dense right central opacity measuring about 6 cm x 3 cm”; the framework would produce the queries “measure the distance between the endotracheal tube and the carina" and “measure the dimensions of the right central opacity”. For this purpose, we utilize GPT-4o mini [14] as the query generator, and details on the prompt used can be found in the Supplementary Material. These queries are then passed on to the Code Generator module.
Code Generator. This module receives the measurement query and a high-level API description. It dynamically generates Python code to answer the query based on the methods provided by the API. We have designed a comprehensive toolbox for identifying and correcting measurement hallucinations in radiology reports, as detailed below:
@@
Report Updater. This module takes the original report together with all the results of the measurement queries to update the report. Specifically, it updates a sentence if a new value is presented, or removes a phrase entirely if the object was hallucinated. Moreover, since the old interpretation may be outdated in light of new evidence, the module references measurement guidelines for each result to determine if the ETT placement should be updated. The pipeline concludes with returning the updated report. For instance, in the case of ETT placement, where the ideal position is 5 ± 2 cm above the carina  [10], if the updated measurement falls outside this range, the Report Updater will flag this discrepancy and update the report to indicate the need for repositioning. We notice that the numeric reasoning of GPT-4o mini [14] is currently insufficient to reliably extract the inference. For instance, the language model fails to identify if a measurement is below or above a threshold value. Therefore, in this version of the pipeline, we implement a strict rule-based approach to infer the placement category of the endotracheal tube. If the distance is between 3-7 centimeters, the placement is deemed as “correct”. Otherwise, the placement is updated to “incorrect”.
@@ 4.2 Tool Modules
  • find(“carina”, “endotracheal tube”): We fine-tune CarinaNet [29] on a diverse private dataset from 22 hospitals. The CarinaNet module, which we call CarinaNet+, outputs an (x, y)-coordinate for both the carina and the endotracheal tube along with confidence scores. See Table 2 for the MIMIC-CXR test set performance.
@@ 4.3 Baselines and Implementation Details
Baseline Models. We demonstrate FactCheXcker’s effectiveness across 11 different medical report generation models, including CheXagent [5], CheXpertPlus [3], GPT-4V [45], LLM-CXR [19], MAIRA-2 [2], MedVersa [48], RadFM [42], RaDialog [31], RGRG [39], VLCI [4]. For all models, we follow the official protocols to generate reports on the MIMIC-CXR test set. Detailed descriptions and implementation specifications are provided in the Supplementary Material.
Measurement Extraction. We implement a systematic extraction process to extract labels of the concrete measurements made in the ground truth- and model-generated reports. To optimize API usage, we first filter reports containing specific measurements using keywords: “cm”, “mm”, “centimeter(s)”, “millimeter(s)”, “measure(s)”. We then utilize GPT-4o mini to extract object names with associated concrete measurements from both ground truth and model-generated reports. The detailed extraction prompt is provided in the Supplementary Material. We categorize the identified measurements into five main categories: Endotracheal tube, Other tubes/catheters (e.g., PICC, central venous line), Lesion (e.g., opacity, mass, nodule), Pneumothorax, and Others (e.g., calcification, balloon pump). Notably, endotracheal tubes account for the most significant proportion of measurements, with an average mention percentage of 78.3% across all models.
Endotracheal Tube Annotation. After extraction of the measurements from the MIMIC-test set, there remained 45 reports where the ground truth report mentioned an endotracheal tube as present, but did not give a concrete measurement. To give more opportunity for comparisons, a trained radiologist annotated the location of the carina and the endotracheal tube tip in the 45 images. Then, we computed the distance between the centers of the annotated bounding boxes as the distance between the carina and the endotracheal tube. This result was passed to the Report Updater-module of our framework, and thus the result was “injected” into the ground truth reports for later comparisons.
@@
ETT Presense ETT Measurement ETT Placement
@@
CheXagent 0.70 0.76 1.98 0.68 191.0% 4.12 1.39 196.0% 0.85 0.90
CheXpertPlus 0.64 0.68 1.44 0.66 118.0% 2.09 0.92 127.0% 0.90 0.94
Cvt2distilgpt2 0.71 0.72 3.82 0.89 329.0% 6.37 1.46 336.0% 0.73 0.88
GPT4V 0.20 0.58 2.35 0.39 503.0% 11.19 1.22 817.0% 1.00 1.00
LLM-CXR 0.08 0.58 2.41 0.91 165.0% 18.54 1.65 1024.0% 0.74 0.97
MAIRA-2 0.63 0.68 1.43 0.99 44.0% 2.01 1.34 50.0% 0.76 0.91
MedVersa 0.73 0.80 1.07 0.86 24.0% 1.49 1.13 32.0% 0.84 0.94
RGRG 0.50 0.60 1.58 0.76 108.0% 2.51 1.09 130.0% 0.77 0.96
RaDialog 0.67 0.69 0.99 0.86 15.0% 1.43 1.23 16.0% 0.81 0.92
RadFM 0.28 0.47 0.65 1.04 -38.0% 8.12 11.56 -30.0% 1.00 1.00
VLCI 0.25 0.54 3.50 0.98 257.0% 21.88 4.90 347.0% 0.80 0.88
Average 0.49 0.65 1.93 0.82 135.0% 7.25 2.54 186.0% 0.84 0.94